OpenStudy (anonymous):

I need help with a graphing quadratic function problem. (please don't solve) I'm stuck on a step. \[f(x)=2x ^{2}-5x+2\] i get this \[2(x ^{2}-\frac{ 5 }{ 4}x+2)\]. Now I know you are supposed to divde b by 2a then square which is 25/16 so i get \[2(x ^{2}-\frac{ 5 }{ 2 }x+\frac{ 25 }{ 16 })=-2+\frac{ 25 }{ 16 }\]

OpenStudy (anonymous):

however, the math lab shows me the -2 being to \[\frac{ 25 }{ 8 }\], where does 25/8 come from, shouldnt it be the same 25\16 for both, or ami missing something

OpenStudy (anonymous):

math lab shows this

OpenStudy (anonymous):

\[2(x ^{2}-\frac{ 5 }{ 2 }x+\frac{ 25 }{ 16 })+2-\frac{ 25 }{ 8 }\]

OpenStudy (anonymous):

If you need to graph this, it's much easier to first compute the intercepts by using the quadratic formula and then you find the y-intercept as well, and just connect the three points to form the parabola.

OpenStudy (anonymous):

\[-b+_\sqrt{b ^{2}}-4(a)c\] over 2(a) , just use this instead of the method im using?

OpenStudy (anonymous):

I would love to know where the 16\8 comes from

OpenStudy (anonymous):

Yeah, you want me to solve this for you so you can apply the method to graphing the other equations u have?

OpenStudy (anonymous):

well the only problem is on the test, i won't be able to use the quadratic function, we have to be able to put the function into this form: \[f(x)=a(x-h)^{2}+k\]

OpenStudy (anonymous):

that is why i'm a little confused on where \[\frac{ 16 }{ 8 }\] comes from

OpenStudy (precal):

ok looks like you are completing the square